# Identifying the post-pandemic determinants of low performing students in Latin America through interpretable Machine Learning SHAP Values—Insights from PISA 2022

### v1

@@ Abstract
The high prevalence of students not achieving the basic competencies in Latin America is concerning. Even more so given the region’s deep structural inequalities and the larger post-pandemic regional learning losses. Within this scenario, this paper contributes to the identification of the determinants of bottom and low performers (below level 2) using recent advancements on explainable machine learning methods. In particular, relying on PISA 2022 data for 10 countries and using the Shapley Additive Explanations (SHAP) analysis, I identify critical factors impacting on the student performance across low performers groups. I find that a student with the highest probability of being a not achiever speaks a minority language and had repeated, has no digital devices at home, comes from a poor family and works for payment half of the week, and the school he//she attends has wide disadvantages such as bad school climate, weak ICT infrastructure and poor teaching quality (only a third of teachers being certified). Regarding countries’ estimates, I find quite homogeneous patterns as far as global average contribution of top ranked factors is concerned, with repetition at primary, household wealth, and educational ICT inputs being top ten ranked covariates in at least 8 out of the 10 total countries. The paper findings contribute to the broad literature on strategies to identify and to target those most left behind in Latin American education systems.
Keywords learning poverty $\cdot$ interpretable machine learning $\cdot$ SHAP values $\cdot$ SDG4 $\cdot$ Latin America $\cdot$ PISA 2022 $\cdot$ COVID-19 pandemic
@@ 1 Introduction
Latin American students’ poor achievement at secondary level (Arias Ortiz et al., 2024b) and weak learning growth (Angrist et al., 2021) are of great policy concern, even more so because the region’s low learning trap is rooted in systemic and profound inequalities (Breton and Canavire-Bacarreza, 2018; Fernández et al., 2024). Based on the OECD Programme for International Student Assessment (PISA) 2022 data, Arias Ortiz et al. (2023) estimate that 75% of the students in the region are low math performers —not reaching the basic competencies given by level 2—, and more than half of students (around 55%) lacks basic reading skills (fail to understand a simple text). Other recent regional learning survey results such the Regional Comparative and Explanatory Study (ERCE 2019) also points in the same direction on the extent of the region’s learning crisis for primary school students (UNICEF et al., 2021).
@@
This educational crisis has been aggravated post-pandemic (Azevedo et al., 2021; Betthäuser et al., 2023; Bracco et al., 2025; Neidhöfer et al., 2021) because the Latin America (henceforth: LAC) region experienced the longest period of school closures with poor connectivity (Gómez and Andrés Uzín P, 2022; Munguia, 2023), where, on average (between 2020 and 2021) schools were closed for 269 days; representing a disruption of 1.42 years (Bracco et al., 2025). Hence, the pandemic pushed a large cohort of students down in the learning distribution affecting the more disadvantaged, with estimates suggesting the LAC region lost around 0.9-1.1 years of schooling (Azevedo et al., 2021). In the case of OECD countries the share of disadvantaged low performers rose by 9% in the 2018-2022 period (OECD, 2023c), whereas in Latin America (already suffering from long term broad educational inequalities Brunori et al., 2024; De la Torre et al., 2017; Delprato et al., 2015) according to Delprato and Antequera (2025), using PISA 2022, the proportion of low performers (below level 2) rose by 5% in math and slightly declined for reading, and remained stable in science (Arias Ortiz et al., 2023). Yet, good cognitive skills are fundamental for quality jobs and to enhance future labour prospects of secondary students which are vital for a steady growth of the economies in the region (Angrist et al., 2021; Hanushek and Woessmann, 2023). It is within this post-pandemic scenario, that the paper addresses the drivers of learning poverty in the LAC region.
@@
The COVID-19 pandemic worsened the already prevailing and detrimental educational inequality in the LAC region. Prolonged school closures led to loss of learning, increased inequalities, having negative impacts on psychosocial health and well-being of students and teachers, with younger children and students from more vulnerable groups experiencing the greatest losses (Asadullah et al., 2023; UNESCO, 2024). For instance, Bracco et al. (2025)’s simulations show how educational loss is tightly related to households’ socioeconomic levels. The authors estimate a mean loss years of education for a student in the bottom decile of the income distribution of 81%, nearly four as large than a student in the top decile (losing 22%).
@@
Household’s income shocks leading to increasing poverty (which in the region are estimated between 23%-39%; Bracco et al., 2025) due to the pandemic and their consequences on educational outputs compounded established drivers of learning poverty in LAC region. Established drivers of low performance in the region are repetition (Randall and Anderson, 2016; Vinas-Forcade et al., 2021), parental education and Indigeneity (Cortina, 2017; Delprato, 2019), educational inputs (e.g., households’ ICT and books stocks, family educational support: Lara and Saracostti, 2019; Torrecilla and Hernández-Castilla, 2020), as well as school characteristics such as school contextual poverty, school type and poor school climate (Berkowitz et al., 2017), ICT investments (number of PCs per student, rate of connectivity; Agasisti et al., 2023), quality of teaching —proxied by teachers uptake of professional development//PD and qualifications (McGinn and Schiefelbein, 2023).
Yet, these drivers are not likely to be homogenous across sub-groups at the bottom of students’ learning distribution, and there is a need of carefully distinguishing these groups for targeting. So, in this paper, I fill this gap by investigating how these determinants may vary along low performing groups in the region. Importantly, this empirical contribution is carried out using the latest PISA available data for the LAC region after the COVID-19 pandemic, offering the latest results on students’ low learning levels determinants. Furthermore, studies using interpretable machine learning (ML) methods based on learning surveys addressing determinants of students’ achievement are rather scarce (e.g., Bu and Chen, 2023; Gómez-Talal et al., 2024; Huang et al., 2024; Liu et al., 2022), with their focus being on richer countries education systems whose realities are entirely contrasting to global south education systems. Methodologically, I contribute here to this small body of literature, expanding it to global south education systems, and at a regional scale.
@@
In particular, relying on the latest post pandemic (year 2022) PISA data for 10 LAC countries, and recent developments in the field of explainable machine learning: Shapley Additive Explanations (SHAP) analysis (Chen et al., 2023; Lundberg and Lee, 2017), this paper attempts to answer the following research questions for students at the bottom of the learning distribution (from level 0 to level 2 achievement groups):
@@
    What are the learning profiles (namely: leading determinants) of students with the highest and lowest chances of poor achievement across the region? This question’s approach has a local dimension as it deals with specific students (at extremes).
@@
    Which are the main drivers at the country level, both globally and locally, of learning poverty?
@@
The rest of the paper is organised as follows. Section 2 contains a review of ML explanatory studies within the educational research field. In Section 3 I describe the data used, the definition of the working samples for the two educational outcomes and leading covariates used. Section 4 describes the methodological approach —i.e., the array of learners (models) and Shap values methodology for interpretability of results and the analytical steps followed. Findings are presented in Section 5, both at the regional level and country level. In Section 6 I discuss the implication of results for future targeting of low performing students in the region and, in Section 7, I include some concluding remarks.
2 Background—machine learning and educational studies
@@
Underpinned by their complexity, machine learning (ML) models are now widely applied to solve supervised learning problems in many different fields and settings (Olsen et al., 2024; Chen et al., 2023; Dhal and Azad, 2022; Sarker, 2021) because supervised ML methods are capable of generating more precise predictions than standard statistical models (Chen et al., 2022; Lisboa et al., 2023). Yet, this may come at the expense of model interpretability (Ruding, 2019). Therefore, to account for this absence of interpretation (Olsen et al., 2024), the field of explainable interpretable machine learning has become an area of recent substantial research (Covert et al., 2021; Linardatos et al., 2020; Holzinger et al., 2020; Molnar, 2020).
@@
In tandem with this methodological progress, there has also been a recent growth the application of ML, broadly in the social sciences (Brand et al., 2023; Grimmer et al., 2021) and within the educational field, which is still limited, but continuously grows (e.g., Ersozlu et al., 2024; Hilbert et al., 2021). Grimmer et al. (2021) advocate for using an agnostic approach, using machine learning to make social science inferences where new research questions are generated, and concepts and hypothesis can be tested with new data, rather than assuming theory are robust before investigating what is emerging from available data. Brand et al. (2023) describe how to use ML methods related to causality, mostly interest in learning about heterogeneous treatment effects and to minimize biases in estimating population-level quantities of interest (also see the review of Athey and Imbens, 2019 for the economics and econometrics fields).
@@
Hilbert et al. (2021) show how the robustness of educational research can benefit from ML methods and the actual prediction of new data and validation of empirical models, specially when the size of data is large like in international learning surveys (e.g., TIMSS, PIRLS, PISA). In the educational research field, so far, most of the focus has been on prediction, such as academic performance in regression tasks (Hellas et al., 2018; Öz et al., 2024) and, to a lesser extent, for the identification of students at risk (e.g., Alam and Mohanty, 2022; Rodríguez et al., 2023). Dol and Jawandhiya (2023), in a review of nearly 150 articles over 2010-2020 period, find out that classification techniques are mostly used technique for analysing students performance within the Educational data mining (EDM) field. Similarly, Ersozlu et al. (2024), in a systematic review of ML methods used for educational data, find that nearly 90% of publications use ML approaches for predicting students’ performances and finding learning patterns. In spite of the prediction-centred approach of ML, researchers have recently produced various approaches to better understanding of models by quantifying the impact and interactions of features (Hilbert et al., 2021). The SHAP method (Lundberg and Lee, 2017), which I employ in the current paper, is a leading example of this, falling into the group of interpretable machine learning (IML) techniques.
@@
Yet, the majority of educational empirical work (so far, a small body of studies) using the SHAP method are either for specific richer countries or regions, without a scaffolding inductive logic of analysis following layers of education systems (Codiroli Mcmaster and Cook, 2019) and how educational inequalities are produced, specially when applied to monitor SDG4 inequalities (Delprato, 2024).
@@
Studies using IML methods based on learning surveys addressing determinants of students’ achievement are rather scanty (Bu and Chen, 2023; Gómez-Talal et al., 2024; Huang et al., 2024; Liu et al., 2022; Zhu et al., 2025), and subject specific. These studies’ focus is on more developed countries with more resources-intensive education systems (e.g., Spain: Gómez-Talal et al., 2024, or the US: Zhu et al., 2025) or groups of countries//regions (East Asia: Bu and Chen, 2023; Huang et al., 2024) whose realities are entirely contrasting to global south education systems. The only exception is Ali et al. (2025), which analyse factors influencing academic performance among primary school students in Somaliland; though, this study, is a combination of ML and regression methods, and with the latter replacing IML techniques as Shapley values. Methodologically, I contribute here to this emerging and scanty group of studies, but expanding it to global south education systems of Latin America as well as at a regional scale.

### v2

# Identifying the post-pandemic determinants of low performing students in Latin America through Interpretable Machine Learning methods

@@ Abstract
Introduction. The high prevalence of students not achieving basic learning competencies in Latin America (LAC) is concerning, even more so considering the region’s deep structural inequalities and the larger post-pandemic learning losses. Within this scenario, the paper aims to contribute to the identification of the determinants of bottom and low performers (below level 2).
Methodology. Based on 2022 data from the Programme for International Student Assessment (PISA) for 10 LAC countries, and using a stacking model integrating binary classification models as well as by applying Shapley Additive Explanations (SHAP) analysis for interpretability, we identify critical factors impacting on the student performance across low performers groups.
Results. We find that a student with the highest probability of being a not achiever speaks a minority language and had repeated, has no digital devices at home, comes from a poor family and works for payment half of the week, and the school the student attends has wide disadvantages such as bad school climate, weak Information and Communication Technology (ICT) infrastructure and poor teaching quality (only a third of teachers being certified). For countries’ estimates, we find quite homogeneous patterns regarding the contribution of top ranked factors, with repetition at primary, household wealth, and educational ICT inputs being top ten ranked covariates in at least 8 out of the 10 total countries.
Discussions. The paper findings contribute to the broad literature on strategies to identify and to target those most left behind in Latin American education systems.
Keywords Learning poverty $\cdot$ Latin America $\cdot$ Programme for International Student Assessment 2022 $\cdot$ COVID-19 pandemic $\cdot$ Interpretable machine learning $\cdot$ Shapley Additive Explanations
@@ 1 Introduction
Latin American students’ poor achievement at secondary level (Arias Ortiz et al., 2024b) and weak learning growth (Angrist et al., 2021) are of great policy concern, even more so because the region’s low learning trap is rooted in systemic and profound inequalities (Breton and Canavire-Bacarreza, 2018; Fernández et al., 2024). Based on the OECD Programme for International Student Assessment (PISA) 2022 data, Arias Ortiz et al. (2023) estimate that 75% of the students in the region are low math performers–not reaching the basic competencies given by level 2–, and more than half of students (around 55%) lacks basic reading skills (fail to understand a simple text). Other recent regional learning survey results such the Regional Comparative and Explanatory Study (ERCE 2019) also points in the same direction on the extent of the region’s learning crisis for primary school students (UNICEF et al., 2021).
@@
This educational crisis has been aggravated post-pandemic (Azevedo et al., 2021; Betthäuser et al., 2023; Bracco et al., 2025; Neidhöfer et al., 2021) because the Latin America (henceforth: LAC) region experienced the longest period of school closures with poor connectivity (Gómez and Andrés Uzín P, 2022; Munguia, 2023) where, on average between 2020 and 2021, schools were closed for 269 days, which represents a disruption of 1.42 years (Bracco et al., 2025). Hence, the pandemic pushed a large cohort of students down in the learning distribution affecting the more disadvantaged, with estimates suggesting the LAC region lost around 0.9-1.1 years of schooling (Azevedo et al., 2021). On top of already suffering from preexistent long term and deep educational inequalities (Brunori et al., 2024; Delprato et al., 2015; Krüger, 2019; Murillo et al., 2023; Neidhöfer et al., 2018), the LAC region experienced an increase on the share of disadvantaged low performers (achieving below level 2) by 5% in math, a slight decline for reading (see: Delprato and Antequera, 2025 using PISA 2022), and with this share being stable for science (Arias Ortiz et al., 2023). Yet, good cognitive skills are fundamental for quality jobs and to enhance future labour prospects of secondary students which are vital for a steady growth of economies in the region (Angrist et al., 2021; Hanushek and Woessmann, 2023). It is within this post-pandemic scenario that the paper addresses the drivers of learning poverty in the LAC region.
@@
In countries where education systems where resources-constrained the negative effects of the COVID-19 pandemic have been more profound because the wider barriers these countries had to tackle given by the triple burden of schooling deprivation, learning inequality and learning poverty (Asadullah et al., 2023). So, in the case of the LAC countries, the pandemic worsened educational inequalities. Prolonged school closures led to loss of learning, increased inequalities, also having negative impacts on psychosocial health and well-being of students and teachers, with younger children and students from more vulnerable groups experiencing the greatest losses (Asadullah et al., 2023; UNESCO, 2024).
@@
In turn, the COVID19 pandemic’s driven households income shocks yielded increasing poverty (which in the region are estimated between 23%-39%; Bracco et al., 2025) and their consequences for educational outputs compounded established drivers of learning poverty in LAC region. For instance, Bracco et al. (2025)’s simulations show how pandemic educational loss in the LAC region is tightly related to households’ socioeconomic levels. The authors estimate a mean loss years of education for a student in the bottom decile of the income distribution of 81%, nearly four times as large than the one experienced by a student in the top decile (losing 22%). In addition, established drivers of low performance in the region are repetition (Randall and Anderson, 2016; Vinas-Forcade et al., 2021), parental education and Indigeneity (Cortina, 2017; Delprato, 2021, 2019), educational inputs (e.g., households’ ICT and books stocks, family educational support: Lara and Saracostti, 2019; Torrecilla and Hernández-Castilla, 2020), as well as school characteristics such as school contextual poverty, school type and poor school climate (Berkowitz et al., 2017; Delprato et al., 2017), schools’ ICT investments (number of personal computers//PC per student, rate of connectivity; Agasisti et al., 2023; CEPAL, 2020), quality of teaching –proxied by teachers uptake of professional development (PD) and qualifications (McGinn and Schiefelbein, 2023).
These drivers, however, are not likely to be homogenous across sub-groups at the bottom of students’ learning distributions. Considering the Programme for International Student Assessment (PISA) learning brackets (OECD, 2023a, b), when targeting students’ proximal determinants (i.e., elements that need to be in place in order for a child to have a successful learning experience in school; Pritchett, 2025), it is not the same if a student does not reach either level 1 (having insufficient foundational skills) or level 2 (reaching a baseline proficiency level), with the furthest disadvantaged students who perform below level 1 (at level 0) not acquiring foundational skills for most basic problems, struggling with fundamental tasks even following direct instructions. Research on learning determinants for the LAC region (e.g., Alves and Candido, 2020; Borchers and da Cunha, 2025) do not place an emphasis on dissecting these students groups between bottom and low performers, and by following an aggregated approach focusing on the drivers of performance for an average student, they are not able to differentiate and identify which modification of malleable students’ factors may allow, for instance, to a student to move up from an extremely low performance (level 0) to a bordering proficiency level (level 1).
As a result, there is a need to carefully distinguish among low performing groups for targeting, even more so considering they make up for the majority of students population (between 55%-75%; Arias Ortiz et al., 2023). In this paper, we therefore attempt to fill this gap by investigating how these determinants may vary along low performing groups for the whole region and specific countries. Importantly, the paper’s empirical contribution is carried out using the latest PISA available data (year 2022) for the LAC region after the COVID-19 pandemic, offering the latest results on students’ low learning levels determinants.
Moreover, studies using interpretable machine learning (IML) methods based on learning surveys addressing determinants of students’ achievement are rather scarce (e.g., Bu and Chen, 2023; Gómez-Talal et al., 2024; Huang et al., 2024; Liu et al., 2022), with their focus being on richer countries education systems whose realities are entirely contrasting to global south education systems. Methodologically, the current study contributes to this small body of literature, expanding the use of IML to assess what drives students from global south education systems (and at a regional scale) to be more prone to move down towards lower achievement brackets in the learning distribution. The study also contributes to the ML (machine learning) literature of the identification of students at risk of failure or dropout within online environments (see, for instance: Shi et al., 2025, 2024) from a regional LAC perspective.
@@
In particular, relying on post-pandemic PISA data for 10 LAC countries, and using recent developments in the field of explainable machine learning: Shapley Additive Explanations (SHAP) analysis (Chen et al., 2023; Lundberg and Lee, 2017), this paper answers the following research questions for students at the bottom of the learning distribution (from level 0 to level 2 achievement groups):
@@
    What are the learning profiles (namely: leading determinants) of students with the highest and lowest chances of being poor achievers across the region? This question’s approach has a local dimension as it deals with specific students (located at extremes).
@@
    What are the main drivers at the country level, both globally and locally, of learning poverty?
@@
The rest of the paper is organised as follows. Section 2 contains the literature review and in Section 3 we describe the data used, the definition of the working samples for the two educational outcomes and leading covariates employed. Section 4 describes the methodological approach –i.e., the array of learners (models) and Shap values methodology for interpretability of results and the analytical steps followed. Findings are presented in Section 5 and in Section 6 we discuss implications of results for future targeting of low performing students in the region. Section 7 contains concluding remarks.
2 Background
2.1 Machine learning in educational research
@@
Underpinned by their complexity, ML models are now widely applied to solve supervised learning problems in many different fields and settings (Chen et al., 2023; Dhal and Azad, 2022; Olsen et al., 2024; Sarker, 2021) because supervised ML methods are capable of generating more precise predictions than standard statistical models (Chen et al., 2022; Lisboa et al., 2023). Yet, this may come at the expense of model interpretability (Ruding, 2019). Therefore, to account for this absence of interpretation (Olsen et al., 2024), the field of IML has become an area of recent substantial research (Covert et al., 2021; Linardatos et al., 2020; Holzinger et al., 2020; Molnar, 2020).
@@
In tandem with this methodological progress, there has also been a recent growth in the application of ML, broadly in the social sciences (Brand et al., 2023; Grimmer et al., 2021) and within the educational field which, although is still limited, it has been continuously growing (e.g., Elbasi et al., 2025; Ersozlu et al., 2024; Hilbert et al., 2021). Grimmer et al. (2021) advocate for using an agnostic approach, using ML to make social science inferences where new research questions are generated and concepts and hypothesis can be tested with new data, rather than assuming a theory before investigating what is emerging from available data. Brand et al. (2023) describe how to use ML methods related to causality, mostly in relationship to heterogeneous treatment effects and how to minimize biases when estimating population-level quantities of interest.1
@@
Hilbert et al. (2021) show how the robustness of educational research can benefit from ML methods and the actual prediction of new data and validation of empirical models, specially when the size of data is large like in international learning surveys (e.g., TIMSS, PIRLS, PISA). So far in the educational research field, most of the focus has been on prediction, such as predicting academic performance in regression tasks (Hellas et al., 2018; Öz et al., 2024) and, to a lesser extent, for the identification of students at risk (e.g., Alam and Mohanty, 2022; Rodríguez et al., 2023). Dol and Jawandhiya (2023), in a review of nearly 150 articles over 2010-2020 period, find that classification techniques are mostly used for analysing students performance within the educational data mining (EDM) field. Similarly, Ersozlu et al. (2024), in a systematic review of ML methods used for educational data, find that nearly 90% of publications use ML approaches for predicting students’ performance and learning patterns. In spite of the prediction-centred approach of ML education studies, researchers have produced various approaches for a better understanding of models by quantifying the impact and interactions of features (Hilbert et al., 2021).
@@
The SHAP method (Lundberg and Lee, 2017), which is employed in the current paper, is a leading example of this, falling into the group of IML techniques. Yet, the majority of educational empirical work using the SHAP method (so far, a small body of studies) are either for specific richer countries or regions, without a scaffolding inductive analytical approach following layers of education systems (Codiroli Mcmaster and Cook, 2019) and how educational inequalities are produced and entrenched when monitoring inequalities (Delprato, 2024).
2.2 Review of machine learning studies–learning surveys and students at risk
@@
Studies using IML methods based on learning surveys addressing determinants of students’ achievement are rather scanty and subject specific (Bu and Chen, 2023; Gómez-Talal et al., 2024; Huang et al., 2024; Liu et al., 2022; Zhu et al., 2025). These studies are focused in better-off and more resources-intensive education systems from developed countries (e.g., Spain: Gómez-Talal et al., 2024, or the US: Zhu et al., 2025) or groups of countries//regions (East Asia: Bu and Chen, 2023; Huang et al., 2024) whose realities are entirely contrasting to global south education systems. An exception is the study by Ali et al. (2025), which addresses factors influencing academic performance among primary school students in Somaliland; though, this study, is a combination of ML and regression methods. The current study contributes to this emerging group of studies, and expanding the empirical evidence on drivers of learning poverty to global south education systems (i.e., the LAC region).
Our paper, through its assessment of determinants of students who are bottom and low performers with pandemic educational data, is related to ML studies dealing with the identification of at-risk students (i.e., failure, dropouts) and, specifically, considering that PISA 2022 data collection addressed issues related to the COVID-19 pandemic and remote learning, it is linked to the prediction of students at risks within online learning environments. A recent review on this topic is given by Shi et al. (2025), which reviews 161 studies published in the 2014-2024 period using deep learning (DL), classical machine learning (cML) and cML ensemble approaches under alternative definitions of students at risk. In this review, the authors highlight: (i) the growth on the use of ensemble of cML models over the last 5 years, having superior performance over DL when following an ensemble approach (with Gradient Boosting Tree and Random Forest emerging as leading cases) and using smaller data sizes; and (ii) the majority of studies tend to use as drivers for students at risk engagement variables and achievement covariates (more widely employed than demographic factors), with demographics factors such as age, ethnicity, disability, socio-economic status and parental education being present in different studies. Moreover, in a meta-regression study looking into the success of ML techniques in identifying online at-risk students by data type and DL and cML approaches, Shi et al. (2024) also find that cML ensemble models exhibit consistently higher prediction accuracy across all evaluated prediction stages.
Specifically, there are various studies about the identification of students at risk (particularity at Higher Education//HE) during the pandemic, with Shi et al. (2025) indicating that around 32% ML studies are based on online HE courses. For example, Khan et al. (2024), is a close study to ours given its focus on the impact of the COVID-19 pandemic on students achievement (using a small dataset from students enrolled during 2020-21 at a Pakistani university), which proposes a regularised ensemble model made up of random forest models to pin down key factors behind their achievement. The authors find that students’ attendance and interaction, internet connectivity, and studying a mandatory entry course during the restrictions are all leading drivers of students’ university achievement. Likewise, Marín et al. (2025), employ an array of nine ML models (with XGBoost showing a high degree of accuracy) to predict grade performance of students in a Peruvian university (for the first semester in 2025), finding that prior academic variables stands out as key predictors, though also non-academic factors such as family support, stress, and time management having a moderate impact on performance. The study emphasises the need of ML for educational research (to account and address high-dimensionality, complex and non-linear associations within educational processes coupled with importance of psychosocial and well-being factors) and the relevance of SHAP values, thereby arguing that they provide a deeper understanding of the individual impact of each variable on the prediction and actionable insights leading to specific observations.
Other similar papers dealing with online learning through ML models at HE are, for instance, Chytas et al. (2023) (for students attending a Greek university) and Asad et al. (2023) (using data from $\approx$ 12,000 Pakistani HE students), with the latter study relying–as the current paper–on a ensemble model constructed from five ML models. Equally, Karalar et al. (2021)’s study relies on an ensemble model (including quadratic discriminant analysis, decision trees, random forest, extra trees, logistic regression, and artificial neural network) to predict students at risk of academic failure in a Turkish public university (n = 2000) during the COVID-19 pandemic, with the model yielding a prediction accuracy of around 90% and a higher relevance of asynchronous learning activities. In another HE context (i.e., Saudi university), Albreiki et al. (2022) propose an ensemble modelling approach with explainable ML and rule-based model risk flags to identify students who are at risk of low performance and their factors. All in all, these empirical studies show the capability of ML model to capture non-linear associations behind the chances of students falling at risk of failure, with a performing superiority of ensemble models.
Our paper innovates the current state-of-the-art of the existing literature in three main ways. Firstly, it is the first empirical investigation which assesses the determinants of subject-combined proficiency by disaggregating weak students’ performance (or learning poverty) into bottom (below level 1) and low (below level 2) students groups for the LAC region, with previous studies limited to richer education systems (in either Europe or Asia) and for higher achievement brackets at given subjects, mostly looking into determinants of proficiency (above level 2). We achieve this by looking into these determinants from a global and local perspective. Secondly, the empirical evidence refers to the most recent wave of internationally comparative data on students’ achievement for the region (i.e., PISA 2022) intertwined by the global COVID-19 pandemic. Thirdly, the empirical analysis includes students from 10 LAC countries, whereas earlier literature concentrates on single countries. Thus, the motivation of this study is to provide a complete overview of determinants of learning poverty in the region as well as on dissimilar patterns driven by country heterogeneity. Finally, the analysis relies on recent progress on IML offering a granular snapshot of leading determinants through the SHAP method, using an ensemble model for better accuracy.